When one mole of hydrogen, H2, is formed, 131 KJ of energy is absorbed.

Calculate the amount of energy absorbed when 240 dm3 of hydrogen, measured at room temperature and pressure is formed.
1
Expert's answer
2020-12-09T06:34:53-0500

According to Avogadro's law 1 mole of any gas occupies 22.4 L (dm3) at room temperature and pressure.

So find the amount of energy absorbed:

"E = 131 * 240\/22.4 = 1403.6 (kJ)"

Answer: 1403.6 kJ is absorbed when 240 dm3 of hydrogen, measured at room temperature and pressure is formed.